Find the values of the six trigonometric functions of theta:
Function Value: sintheta=3/8
Constraint: costheta<0

Answer by lwsshak3(11628)

Find the values of the six trigonometric functions of theta:
Function Value: sintheta=3/8
Constraint: costheta<0
***
Given data shows reference angle is in quadrant II in which sin>0, cos<0
Use x for theta:
sinx=3/8
cosx=-√1-sin^2(x)=-√(1-9/64)=-√(55/64)=-√55/8
tanx=sinx/cosx=-3/√55=-3√55/55
cscx=8/3
secx=-8/√55=-8√55/55
cotx=-√55/3